brittney12:

An employer uses the linear regression equation y=0.18x+320.22 to predict the weekly salary, y, of an employee who sells x dollars worth of merchandise. Last week, Joaquin sold $1500 worth of merchandise. The same week, Alex earned $650. Using the regression equation, which is an accurate comparison?

dude:

Hmm can you take a screenshot of the graph and post it here?

Champion:

Hello @brittney12 Your question asks "which", which suggests that your question should have multiple choice options. Without knowing these options, it would be difficult for us to help you. In the mean time, I'll try to do what I can to help you through this problem:

Champion:

You are given that y = 0.18x + 320.22 is the equation the employer is using. And that x stands for the "dollars worth of merchandise" sold by each employee. "Joaquin sold $1500 worth of merchandise" This means we should plug in 1500 for x y = 0.18(1500) + 320.22 = 270 + 320.22 = 590.22 so Joaquin's weekly salary is going to be $590.22 notice that "Alex earned $650" is different this phrase is saying that Alex has a weekly salary of $650, NOT that Alex sold $650 worth of merchandise. This means we should plug in 650 for y and NOT x 650 = 0.18x + 320.22 0.18x = 329.78 x = 1832.11 so Alex sold $1832.11 worth of merchandise.

Champion:

This is as much as I can help you with without knowing your multiple choice options. Feel free to come back any time to ask for more help! Good luck :-)